Find the first term of an A.p if the 5th term is 28 and sum of the first 17th term is Zero
Answer by mananth(16946)

Find the first term of an A.p if the 5th term is 28 and sum of the first 17th term is Zero
tn = a+(n-1)d
28 = a+4d
a+4d= 28....................(1)
Sn = n/2( 2a+(n-1)d))
0= (17/2)(2a+16d)
0= 17(2a+16d)
34a+272d=0
/34
a+8d=0.....................(2)
subtract (2) from (1)
we get -4d=28
d= -7
Now a+4d=28
a+4*(-7)=28
a-28=28
a=56